OpenStudy (anonymous):

The population of a town was 57,000 in 1980 and 81,000 in 2010. Assume the rate of population increase is constant. What will the population be in 2015? A. 81,400 B. 81,800 C. 83,000 D. 85,000

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(57000\quad ,&1980)\quad &(81000\quad ,&2010) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\\ \quad \\ \textit{how many more folks will there be in 2015, or 5 years later?}\\ \quad \\ m\cdot 5\quad \textit{in 5 years}\\ \quad \\ \begin{array}{llll} 81000+(&m\cdot 5)\implies \textit{population in 2015}\\ \uparrow &\uparrow \\ 2010+&\textit{5 more years} \end{array}\)